OpenStudy (anonymous):

Connie needs to burn at least 250 calories at the gym today. She can burn 22 calories per minute on the treadmill and 10 calories per minute on the elliptical trainer. Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 250 calories?

OpenStudy (anonymous):

\[22t - 10e \ge 250\] \[22t + 10e \le 250 \] \[22t + 10e \ge 250\] \[22t - 10e \le 250\]